# Sharp thresholds for spanning regular subgraphs

### v2

@@ 6.3. Distribution of fragments
Let $\mathbf{W}\subset{[n]\choose 2}$ be a uniformly random $m$-subset. Let $\mathbf{F}^{\prime}\sim\mathcal{Q}_{F,\mathbf{W}}$. We get that whp $|\mathbf{F}^{\prime}\cap F|\leq\ell_{0}+5\chi$. Moreover, by the definition of this distribution, whp, in the $\mathbf{W}$-measure, for every set $\mathcal{X}$ of disjoint closed subgraphs of $F$ on at least $\sqrt{n}/\ln n$ vertices, the $\mathcal{Q}_{F,\mathbf{W}}$-probability that the set of inclusion-maximal closed subgraphs of $\mathbf{F}^{\prime}\cap F$ on at least $\sqrt{n}/\ln n$ vertices contains $\mathcal{X}$ is at most $\left\lfloor\frac{\ln^{2}n}{n^{\delta}}\right\rfloor$. We call a pair $(F,W)$ with $\mathcal{Q}_{F,W}$ satisfying this probability bound, separating. The property of being separating will be applied in Section 6.7 to show that we may cover at least one fragment of size $o(\log n)$ whp.
@@ 6.4. Reconstruction from a smoothed fragment
Let us fix any $W\in{{[n]\choose 2}\choose m}$. Let $\Sigma_{W}$ be the set of all $F\in\mathcal{F}_{n}$ such that $(F,W)\in\Sigma$ is separating. Due to (6.2) and the conclusion of Section 6.3, whp, for (asymptotically) almost all $F$, we have that $F\in\Sigma_{\mathbf{W}}$.
Now, for every pair $(F,W)$ such that $F\in\Sigma_{W}$, let $\mathbf{F}^{\prime}\sim\mathcal{Q}_{F,W}$ be generated independently (mutually over all such pairs $(F,W)$). Let $H(F):=F\cap\mathbf{F}^{\prime}$, and let $H^{\prime}(F)$ be the smoothed version of $H(F)$ (with respect to the matching rule $B$). For brevity, we denote $\ell_{\chi}:=\ell_{0}+5\chi.$
@@
$|\mathcal{S}(S)|\leq 64^{\ell(S)}(n-x(S)+c(S))!\max\left\{n^{c(S)/6},{\ell_{\chi}\choose c(S)+\sigma(S)}\right\}.$ (38)
@@
$|\mathcal{S}(S)|\leq 64^{\ell(S)}(n-x(S)+c(S))!\frac{\ell_{\chi}^{c(S)+\sigma(S)}}{\min\left\{(\ln n)^{c(S)+\sigma(S)},n^{1/4}\right\}}$ (39)
holds with $\mathcal{Q}$-probability $1-e^{-\Omega(n^{1/6})}$.
@@
For brevity, set $c:=c(S)$, $\ell:=\ell(S)$, $x:=x(S)$, and $\sigma:=\sigma(S)$. Let $T_{1},\ldots,T_{c}$ be the connected components of $S$. We fix an arbitrary root in every connected component of $S$ and order edges in every connected component in a way such that edges that are closer to the root appear earlier. Let
@@
The tuple $(\alpha,A,A^{\prime})$ identifies two non-negative integers $a,a^{\prime}$ such that $a+a^{\prime}\leq\max\{c,\chi\}$ that will be explicitly defined later. We also let
  • $\tau_{1}:[a]\to[\lfloor 2n^{2/3}/w\rfloor]$, $\tau_{2}:[a]\to[\lfloor\ln n/4\rfloor]$, $\tau_{3}:[\sigma]\to[\ell_{\chi}]$ be arbitrary functions;
@@
Let us show that the tuple $\mathbf{x}:=(A,A^{\prime},\alpha,\tau_{1},\tau_{2},\tau_{3},f,\rho,\pi)$ identifies a graph $F\in\Sigma_{W}$ such that $S\subset H^{\prime}(F)$. It would immediately imply the desired assertion. Below, we describe the algorithm of reconstruction of $F$ from $\mathbf{x}$.
@@
    Choose pairs of edges in $S$ that were glued in $H(F)$: For every connected component $T_{i}\subset S$, let $A\cap T_{i}=\{v^{i}_{1},\ldots,v^{i}_{t_{i}}\}$, where $v^{i}_{1}<_{\alpha}\ldots<_{\alpha}v^{i}_{t_{i}}$. We also assume that $v^{i}_{t_{i}}$ is does not equal to the $\alpha$-biggest vertex in $T_{i}$. We treat each vertex $v^{i}_{j}$ as the biggest vertex of $S$ that belongs to the smaller edge of the $j$-th pair of edges that was glued. For $j\in[t_{i}+1]$, let $P^{i}_{j}$ be the subgraph of $T_{i}$ induced by the set of all vertices that are $\alpha$-bigger than $v^{i}_{j-1}$ and do not $\alpha$-exceed $v^{i}_{j}$. Here, $v^{i}_{0}=0$ and $v^{i}_{t_{i}+1}$ is the $\alpha$-biggest vertex of $T_{i}$.
@@
    Let $\mathcal{I}^{\prime}$ be the set of all $i\in[c]$ such that $A^{\prime}\cap T_{i}$ is non-empty and $v^{i}_{t_{i}}$ is not the $\alpha$-biggest vertex of $T_{i}$. We also let $a^{\prime}=\sum_{i\in\mathcal{I}^{\prime}}|A^{\prime}\cap T_{i}|$, i.e. $a^{\prime}=\sum_{i}t_{i}$. For every $i\in\mathcal{I}^{\prime}$ and every $j\in[t_{i}]$, we choose the distance between the biggest vertex in $P^{i}_{j}$ and the smallest vertex in $P^{i}_{j+1}$, according to $f$. Note that $f$ is indeed sufficient: since pairs of different glued edges do not belong to the same inclusion-maximal closed subgraph of the fragment $H(F)$, we have that every $v_{i}^{j}$, $j\geq 2$, contributes at least one missing edge in $\ell$ to $\sigma$. Thus, $f$ identifies the position of every $P^{i}_{2},\ldots,P^{i}_{t_{i}+1}$, as soon as the position of $P^{i}_{1}$ is fixed.
@@
    We note that, according to the definition of a smoothed fragment, there is no components in $S$ that contains deleted pieces and glued pairs of edges simultaneously. It means that $\mathcal{I}\cap\mathcal{I}^{\prime}=\varnothing.$ Therefore, it remains to order all the remaining $n-x+c-a$ connected components of $S$ (including the remaining singletons) according to the permutation $\pi$ and complete the reconstruction of $F$.
@@
$\max_{a\leq\max\{c,\chi\},a+a^{\prime}\leq c+\sigma}$ $4^{\ell}2^{2x}\left(\frac{n^{2/3}\ln n}{2w}\right)^{a}\frac{\chi!}{(\chi-a)!}{\ell_{\chi}\choose a^{\prime}}(n-x+c-a)!$
$\leq 64^{\ell}(n-x+c)!\max_{q\leq c,a+a^{\prime}\leq c+\sigma}\left(\frac{\chi n^{2/3}\ln n/(2w)}{n-x+c-a}\right)^{a}{\ell_{\chi}\choose a^{\prime}}$
$\leq 64^{\ell}(n-x+c)!\max\left\{\left(\frac{\chi n^{2/3}\ln n/(2w)}{n-x}\right)^{c},{\ell_{\chi}\choose c+\sigma}\right\}.$
@@
[…]
$\leq 64^{\ell}(n-x+c)!\max\left\{n^{c/6},{\ell_{\chi}\choose c+\sigma}\right\}.$
Since $n^{c/6}<\frac{\ell_{\chi}^{c+\sigma}}{(\ln n)^{c+\sigma}+n^{1/4}}$, it remains to prove that with sufficiently large probability the factor (${\ell_{\chi}\choose c+\sigma}$ can be replaced with $\frac{\ell_{\chi}^{c+\sigma}}{\min\{(\ln n)^{c+\sigma},n^{1/4}\}}$. This makes a significant difference when $c+\sigma=o(\log n/\log\log n)$. Let us recall that the factor (${\ell_{\chi}\choose c+\sigma}$ appears due to the choice of the partition $f$ in step (5) of the reconstruction procedure. If we skip this particular step, then, after implementing all the other steps, we get that $F$ is reconstructed, up to the choice of the tuple $(q_{1},\ldots,q_{a^{\prime}})$ of lengths (i.e., the number of vertices) of closed subgraphs that were cut from $H(F)$. For each $q_{i}$, we first decide, whether $q_{i}\leq\sqrt{n}/\ln n$. There are at most $2^{c(S)+\sigma(S)}$ ways to make this choice. For each $q_{i}$ that should not exceed $\sqrt{n}/\ln n$, we simply choose its value in at most $\sqrt{n}/\ln n<\ell_{\chi}/2$ ways. Without loss of generality, let $q_{1},\ldots,q_{a^{\prime\prime}}$ be the lengths exceeding the threshold $\sqrt{n}/\ln n$. Each evaluation of these lengths identifies an $F\in\mathcal{F}_{n}$. We only count evaluations such that $F\in\Sigma_{W}$. In particular, it means that $(F,W)$ is separating, and thus the probability that $H(F)$ contains a specific subset of inclusion-maximal closed subgraphs — defined by the specified $q_{1},\ldots,q_{a^{\prime\prime}}$ — is at most $n^{-1/3}$. The random variable $\eta$ that counts the number of evaluations of $q_{1},\ldots,q_{a^{\prime\prime}}$ such that $H(F=F(q_{1},\ldots,q_{a^{\prime\prime}}))$ contains the respective subset of inclusion-maximal closed subgraphs is stochastically dominated by $\mathrm{Bin}\left((\ell_{\chi}/2)^{a^{\prime\prime}},n^{-1/3}\right)$, where $(\ell_{\chi}/2)^{a^{\prime\prime}}>\sqrt{n}$. Therefore,
$\mathbb{P}\left(\eta>\frac{(\ell_{\chi}/2)^{a^{\prime\prime}}}{n^{1/4}}\right)=e^{-\Omega(n^{1/6})}.$
@@
We are now ready for the second fragmentation round. From the previous round, we have a sufficiently large multiset of smoothed fragments. Due to Claim 6.7, we can evaluate all $\mathbf{F}^{\prime}\sim\mathcal{Q}_{F,W}$, for every pair $(F,W)$ satisfying $F\in\Sigma_{W}$, in a way such that (38) holds for all $S$ and (39) holds for all $S$ with $\ell(S)\leq n^{0.1}$. The latter inequality is due to the union bound over all such $S$ and the probability bound from (39). We fix such an evaluation $F^{\prime}=F^{\prime}(F,W)$ for every pair $(F,W)$ satisfying $F\in\Sigma_{W}$. Then, for every $F\in\mathcal{F}_{n}$, such that $F\in\Sigma_{\mathbf{W}}$, we include the smoothed version of $F\cap F^{\prime}$ into the new multiset $\mathcal{H}^{(1)}$. Recall that $|\mathcal{H}^{(1)}|$ equals the number of all $F$ such that $F\in\Sigma_{\mathbf{W}}$, which is $(1-o(1))n!$ whp. Clearly, we may assume that every member of this family has size exactly $\ell_{0}+5\chi$.
@@
$p^{H}(\ell,x,c)\leq\frac{|\mathcal{J}^{H}_{\ell,x,c}|\cdot 64^{\ell}(n-x+c)!\max\{n^{c/6},{\ell_{\chi}\choose c+\sigma}\}}{|\mathcal{H}^{(1)}|}.$
@@
$\Pi^{H}_{\ell}=\sum_{c=1}^{\ell}\sum_{\ell/2+3c/2\leq x\leq\ell+c}p_{\ell,x,c}\leq\sum_{c,x}\frac{{\ell_{\chi}\choose c}2^{12\ell}\left(n-x+c\right)!\max\{n^{c/6},{\ell_{\chi}\choose c+\sigma}\}}{(1-o(1))n!}.$
@@
$\Pi^{H}_{\ell}=O\left(\sum_{c,x}\frac{{\ell_{\chi}\choose c}2^{12\ell}\max\{n^{c/6},{\ell_{\chi}\choose c+\sigma}\}}{n^{\ell/2+c/2+\sigma/2}}\right).$
@@
This implies
($\frac{{\ell_{\chi}\choose c}2^{12\ell}n^{c/6}}{n^{\ell/2+c/2+\sigma/2}}\left((1+o(1))\frac{N}{m^{\prime}}\right)^{\ell}$ $\leq\frac{\left(\frac{e\ell_{\chi}}{c}\right)^{c}(2^{12}/\varepsilon+o(1))^{\ell}}{n^{c/3+\sigma/2}}$
$\leq\left(\frac{3Cn^{1/2-1/3+1/50}}{c}\right)^{c}\left(\frac{2^{12}}{\varepsilon}+o(1)\right)^{\ell}n^{-2C\ell/\ln n}$
$<\left(\frac{n^{0.2}}{c}\right)^{c}e^{-C\ell}$
if $C=C(\varepsilon)$ is large enough. In the case when maximum is achieved at (${\ell_{\chi}\choose c+\sigma}$, we can clearly assume that $c+\sigma<\ell_{\chi}/2$ since otherwise
($\frac{{\ell_{\chi}\choose c}2^{12\ell}{\ell_{\chi}\choose c+\sigma}}{n^{\ell/2+c/2+\sigma/2}}\left((1+o(1))\frac{N}{m^{\prime}}\right)^{\ell}=n^{-\Omega(\sqrt{n})}.$
With this assumption, we get that (${\ell_{\chi}\choose c}\leq{\ell_{\chi}\choose c+\sigma}$. Therefore,
($\frac{{\ell_{\chi}\choose c}2^{12\ell}{\ell_{\chi}\choose c+\sigma}}{n^{\ell/2+c/2+\sigma/2}}\left((1+o(1))\frac{N}{m^{\prime}}\right)^{\ell}$ $\leq\frac{\left(\frac{e\ell_{\chi}}{c+\sigma}\right)^{2(c+\sigma)}(2^{12}/\varepsilon+o(1))^{\ell}}{n^{c/2+\sigma/2}}$
$\leq\left(\frac{9C^{2}n^{1/2}}{(c+\sigma)^{2}}\right)^{c+\sigma}\left(\frac{2^{12}}{\varepsilon}+o(1)\right)^{\frac{\sigma\ln n}{4C}+\frac{c\ln n}{100C}}<\left(\frac{n^{0.55}}{(c+\sigma)^{2}}\right)^{c+\sigma}.$

### v3

@@ 6.3. Distribution of fragments
    for every set $\mathcal{X}$ of disjoint closed subgraphs of $F$ on at least $\sqrt{n}/\ln n$ vertices, the $\mathcal{Q}_{F,\mathbf{W}}$-probability that the set of inclusion-maximal closed subgraphs of $\mathbf{F}^{*}\cap F$ on at least $\sqrt{n}/\ln n$ vertices contains $\mathcal{X}$ is at most $\left\lfloor\frac{\ln^{2}n}{n^{\delta}}\right\rfloor$.
We call a pair $(F,W)$ with $\mathcal{Q}_{F,W}$ satisfying P3 separating. The property of being separating will be applied in Section 6.7 to show that we may cover at least one fragment of size $o(\log n)$ whp.
@@ 6.4. Reconstruction from a smoothed fragment
For every pair $\left(F\in\mathcal{F}_{n},W\in{{[n]\choose 2}\choose m}\right)$, let $\mathbf{F}^{\prime}\sim\mathcal{Q}_{F,W}$ be generated independently (mutually over all such pairs $(F,W)$). Recall the definition of $\Sigma$ from Section 6.2. For $W\in{{[n]\choose 2}\choose m}$, let $\Sigma_{W}$ be the set of all $F\in\mathcal{F}_{n}$ such that $(F,W)\in\Sigma$ is separating. Due to (6.2) and the conclusion of Section 6.3, $(\mathbf{W}\times\mathcal{Q})$-whp, for (asymptotically) almost all $F$, we have that $F\in\Sigma_{\mathbf{W}}$.
We now fix $W\in{{[n]\choose 2}\choose m}$ and, for every $F\in\mathcal{F}_{n}$, generate $\mathbf{F}^{\prime}\sim\mathcal{Q}_{F,W}$. If $F\in\Sigma_{W}$, we let $H(F):=F\cap\mathbf{F}^{\prime}$, and let $H^{\prime}(F)$ be the smoothed version of $H(F)$ (with respect to the matching rule $B$). For brevity, we denote $\ell_{\chi}:=\ell_{0}+5\chi.$
@@
$|\mathcal{S}(S)|\leq 64^{\ell}(n-x+c)!\max\left\{n^{c/6}{\ell_{\chi}\choose\min\{\sigma,\lfloor\ell_{\chi}/2\rfloor\}},{\ell_{\chi}\choose\min\{c+\sigma,\lfloor\ell_{\chi}/2\rfloor\}}\right\}\quad\text{ almost surely}.$ (43)
@@
$|\mathcal{S}(S)|\leq 64^{\ell}(n-x+c)!\frac{(2\ell_{\chi})^{c+\sigma}}{\min\left\{(\ln n)^{c+\sigma},n^{1/4}\right\}}$ (44)
holds with $\mathcal{Q}$-probability $1-e^{-\Omega(n^{1/6})}$ (recall that $\mathcal{Q}=\times\mathcal{Q}_{F,W}$).
@@
Let $T_{1},\ldots,T_{c}$ be the connected components of $S$. We fix an arbitrary root in every connected component of $S$ and order edges in every connected component in a way such that edges that are closer to the root appear earlier. Let
@@
The tuple $(\alpha,A,A^{\prime})$ identifies two non-negative integers $a,a^{\prime}$ such that $a\leq\min\{c,\chi\}$ and $a+a^{\prime}\leq c+\sigma$ that will be explicitly defined later. We also let
  • $\tau_{1}:[a]\to[\lfloor 2n^{2/3}/w\rfloor]$, $\tau_{2}:[a]\to[\lfloor\ln n/4\rfloor]$ be arbitrary functions;
@@
Let us show that the tuple $\mathbf{x}:=(A,A^{\prime},\alpha,\tau_{1},\tau_{2},f,\rho,\pi)$ identifies a graph $F\in\Sigma_{W}$ such that $S\subset H^{\prime}(F)$. As we will see, the latter implies the desired assertion. Below, we describe the algorithm of reconstruction of $F$ from $\mathbf{x}$.
@@
    Choose pairs of edges in $S$ that were glued in $H(F)$: For every connected component $T_{i}\subset S$, let $A^{\prime}\cap T_{i}=\{v^{i}_{1},\ldots,v^{i}_{t_{i}}\}$, where $v^{i}_{1}<_{\alpha}\ldots<_{\alpha}v^{i}_{t_{i}}$. We also assume that $v^{i}_{t_{i}}$ does not equal to the $\alpha$-biggest vertex in $T_{i}$. We treat each vertex $v^{i}_{j}$ as the biggest vertex of $S$ that belongs to the smaller edge of the $j$-th pair of edges that was glued. For $j\in[t_{i}+1]$, let $P^{i}_{j}$ be the subgraph of $T_{i}$ induced by the set of all vertices that are $\alpha$-bigger than $v^{i}_{j-1}$ and do not $\alpha$-exceed $v^{i}_{j}$. Here, $v^{i}_{0}=0$ and $v^{i}_{t_{i}+1}$ is the $\alpha$-biggest vertex of $T_{i}$.
@@
    Let $\mathcal{I}^{\prime}$ be the set of all $i\in[c]$ such that $A^{\prime}\cap T_{i}$ is non-empty and $v^{i}_{t_{i}}$ is not the $\alpha$-biggest vertex of $T_{i}$. We also let $a^{\prime}=\sum_{i\in\mathcal{I}^{\prime}}|A^{\prime}\cap T_{i}|$, i.e. $a^{\prime}=\sum_{i\in\mathcal{I}^{\prime}}t_{i}$. We note that $a+a^{\prime}\leq c+\sigma$ since 1) the moved pieces and the glued edges do not belong to the same connected component in the smoothed fragment $H^{\prime}(F)$, according to its construction, 2) different pairs of glued edges do not belong to the same inclusion-maximal closed subgraph of the fragment $H(F)$. For every $i\in\mathcal{I}^{\prime}$ and every $j\in[t_{i}]$, we choose the distance between the biggest vertex in $P^{i}_{j}$ and the smallest vertex in $P^{i}_{j+1}$, according to $f$, that identifies the positions of all $P^{i}_{2},\ldots,P^{i}_{t_{i}+1}$, $i\in\mathcal{I}^{\prime}$, as soon as the positions of $P^{i}_{1}$ are fixed.
@@
    Since there is no component in $S$ that contains deleted pieces and glued pairs of edges simultaneously, we have that $\mathcal{I}\cap\mathcal{I}^{\prime}=\varnothing.$ Therefore, it remains to order all the remaining $n-x+c-a$ connected components of $S$ (including the remaining singletons) according to the permutation $\pi$ and complete the reconstruction of $F$.
@@
$\max_{a\leq\min\{c,\chi\},\,a+a^{\prime}\leq c+\sigma}$ $4^{\ell}2^{2x}\left(\frac{n^{2/3}\ln n}{2w}\right)^{a}\frac{\chi!}{(\chi-a)!}{\ell_{\chi}\choose a^{\prime}}(n-x+c-a)!$
$\leq 64^{\ell}(n-x+c)!\max_{a\leq c,\,a+a^{\prime}\leq c+\sigma}\left(\frac{\chi n^{2/3}\ln n/(2w)}{n-x+c-a}\right)^{a}{\ell_{\chi}\choose a^{\prime}}.$
@@
$\left(\frac{\chi n^{2/3}\ln n/(2w)}{n-x+c-a}\right)^{a}\leq\left(\frac{\chi n^{2/3}\ln n/(2w)}{n-x}\right)^{a}\leq\left(\frac{n^{7/6}}{2(n-2\ell_{0}-4\chi)}\right)^{a}\leq n^{a/6}.$
For a fixed $a+a^{\prime}=:a^{\prime\prime}$, the function $n^{a/6}{\ell_{\chi}\choose a^{\prime\prime}-a}$ is convex in $a$, implying that
$\max_{a\leq c}n^{a/6}{\ell_{\chi}\choose a^{\prime\prime}-a}=\max\left\{n^{c/6}{\ell_{\chi}\choose a^{\prime\prime}-c},{\ell_{\chi}\choose a^{\prime\prime}}\right\}.$
Therefore,
$|\mathcal{S}(S)|\leq 64^{\ell}(n-x+c)!\max\left\{n^{c/6}{\ell_{\chi}\choose\min\{\sigma,\lfloor\ell_{\chi}/2\rfloor\}},{\ell_{\chi}\choose\min\{c+\sigma,\lfloor\ell_{\chi}/2\rfloor\}}\right\}.$
The proof of (43) is completed.
It remains to prove that with sufficiently large probability the factor (${\ell_{\chi}\choose a^{\prime}}$ can be replaced with $\frac{(2\ell_{\chi})^{a^{\prime}}}{\min\{(\ln n)^{a^{\prime}},n^{1/4}\}}$. Indeed, if this is the case, then, since, for a fixed $a+a^{\prime}=:a^{\prime\prime}$, the function $n^{a/6}\frac{\ell_{\chi}^{a^{\prime\prime}-a}}{\min\{(\ln n)^{a^{\prime\prime}-a},n^{1/4}\}}$ decreases in $a$, we get
$\max_{a\leq c}n^{a/6}\frac{(2\ell_{\chi})^{a^{\prime\prime}-a}}{\min\{(\ln n)^{a^{\prime\prime}-a},n^{1/4}\}}=\frac{(2\ell_{\chi})^{a^{\prime\prime}}}{\min\{(\ln n)^{a^{\prime\prime}},n^{1/4}\}}\leq\frac{(2\ell_{\chi})^{c+\sigma}}{\min\{(\ln n)^{c+\sigma},n^{1/4}\}},$
as required in (44).
This refinement makes a significant difference when $c+\sigma=o(\log n/\log\log n)$. Let us recall that the factor (${\ell_{\chi}\choose c+\sigma}$ appears due to the choice of the partition $f$ in step (5) of the reconstruction procedure. If we skip this particular step, then, after implementing all the other steps, we get that $F$ is reconstructed, up to the choice of the tuple $(q_{1},\ldots,q_{a^{\prime}})$ of lengths (i.e., the number of vertices) of closed subgraphs that were cut from $H(F)$. For each $q_{i}$, we first decide, whether $q_{i}\leq\sqrt{n}/\ln n$. There are at most $2^{c(S)+\sigma(S)}$ ways to make this choice. For each $q_{i}$ that should not exceed $\sqrt{n}/\ln n$, we simply choose its value in at most $\sqrt{n}/\ln n<\ell_{\chi}/\ln n$ ways. Without loss of generality, let $q_{1},\ldots,q_{a^{\prime\prime}}$ be the unspecified lengths exceeding the threshold $\sqrt{n}/\ln n$. Each evaluation of these lengths identifies an $F\in\mathcal{F}_{n}$. We only count evaluations such that $F\in\Sigma_{W}$. In particular, it means that $(F,W)$ is separating (i.e. $\mathcal{Q}_{F,W}$ has property P3), and thus, for any $F\in\Sigma_{W}$, the probability that $H(F)$ contains a specific subset of inclusion-maximal closed subgraphs — defined by the specified $q_{1},\ldots,q_{a^{\prime\prime}}$ — is at most $n^{-1/3}$. The random variable $\eta$ that counts the number of evaluations of $q_{1},\ldots,q_{a^{\prime\prime}}$ such that $H(F=F(q_{1},\ldots,q_{a^{\prime\prime}}))$ contains the respective subset of inclusion-maximal closed subgraphs is stochastically dominated by $\mathrm{Bin}\left(\ell_{\chi}^{a^{\prime\prime}},n^{-1/3}\right)$, since $\mathcal{Q}$ is a product measure. Since $\ell_{\chi}^{a^{\prime\prime}}>\sqrt{n}$, we get
$\mathbb{P}\left(\eta>\frac{\ell_{\chi}^{a^{\prime\prime}}}{n^{1/4}}\right)=e^{-\Omega(n^{1/6})}.$
@@
We are now ready for the second fragmentation round. Let us recall that we sampled $\mathbf{F}^{\prime}\sim\mathcal{Q}_{F,W}$ and got a fragment $H(F,W)=F\cap\mathbf{F}^{\prime}$ for every pair $(F,W)$. Then we defined a smoothed version $H^{\prime}(F,W)$ of $H(F,W)$ for every $W$ and every $F\in\Sigma_{W}$. Due to Claim 6.7, (43) holds for all $S$ and (44) holds for all $S$ with $\ell(S)\leq n^{0.1}$, $\mathcal{Q}$-whp. The latter inequality is due to the union bound over all such $S$ and the probability bound from (44). Therefore, we can evaluate all $\mathbf{F}^{\prime}=F^{\prime}$ and treat them as deterministic graphs such that
  • $|\Sigma_{\mathbf{W}}|=(1-o(1))n!$ whp;
  • (43) holds for all $S$;
  • (44) holds for all $S$ with $\ell(S)\leq n^{0.1}$ deterministically.
Let us consider the (random) multiset $\mathcal{H}^{(1)}$ of all $H^{\prime}(F)$ over $F\in\Sigma_{\mathbf{W}}$. Then
$|\mathcal{H}^{(1)}|=|\Sigma_{\mathbf{W}}|=(1-o(1))n!$
whp. Clearly, we may assume that every member of this family has size exactly $\ell_{0}+5\chi$.
@@
$p^{H}(\ell,x,c)\leq\frac{|\mathcal{J}^{H}_{\ell,x,c}|\cdot 64^{\ell}(n-x+c)!\max\left\{n^{c/6}{\ell_{\chi}\choose\min\{\lfloor\ell_{\chi}/2\rfloor,\sigma\}},{\ell_{\chi}\choose\min\{\lfloor\ell_{\chi}/2\rfloor,c+\sigma\}}\right\}}{|\mathcal{H}^{(1)}|}.$
@@
$\Pi^{H}_{\ell}$ $=\sum_{c=1}^{\ell}\sum_{\ell/2+3c/2\leq x\leq\ell+c}p_{\ell,x,c}$
$\leq\sum_{c,x}\frac{{\ell_{\chi}\choose c}2^{12\ell}\left(n-x+c\right)!\max\left\{n^{c/6}{\ell_{\chi}\choose\min\{\lfloor\ell_{\chi}/2\rfloor,\sigma\}},{\ell_{\chi}\choose\min\{\lfloor\ell_{\chi}/2\rfloor,c+\sigma\}}\right\}}{(1-o(1))n!}.$
@@
$\Pi^{H}_{\ell}=O\left(\sum_{c,x}\frac{{\ell_{\chi}\choose c}2^{12\ell}\max\left\{n^{c/6}{\ell_{\chi}\choose\min\{\lfloor\ell_{\chi}/2\rfloor,\sigma\}},{\ell_{\chi}\choose\min\{\lfloor\ell_{\chi}/2\rfloor,c+\sigma\}}\right\}}{n^{\ell/2+c/2+\sigma/2}}\right).$ (47)
@@
Assume first
$n^{c/6}{\ell_{\chi}\choose\min\{\lfloor\ell_{\chi}/2\rfloor,\sigma\}}\geq{\ell_{\chi}\choose\min\{\lfloor\ell_{\chi}/2\rfloor,c+\sigma\}}.$
If, also, $\lfloor\ell_{\chi}/2\rfloor\leq\sigma$, then, due to (48) and (47),
$\Pi^{H}_{\ell}\left((1+o(1))\frac{N}{m^{\prime}}\right)^{\ell}$ ($\leq\frac{{\ell_{\chi}\choose c}2^{12\ell}n^{c/6}2^{2\sigma}}{n^{\ell/2+c/2+\sigma/2}}\left((1+o(1))\frac{N}{m^{\prime}}\right)^{\ell}$ (49)
$\leq\frac{\left(\frac{e\ell_{\chi}}{c}\right)^{c}(2^{12}/\varepsilon+o(1))^{\ell}2^{2\sigma}}{n^{c/3}n^{\sigma/2}}$
$\leq\left(\frac{3Cn^{1/6+1/50}}{c}\right)^{c}\left(\frac{2^{24}}{\varepsilon}+o(1)\right)^{\ell}n^{-2C\ell/\ln n}<\left(\frac{n^{0.2}}{c}\right)^{c}e^{-C\ell}$ (50)
if $C=C(\varepsilon)$ is large enough. If $\lfloor\ell_{\chi}/2\rfloor>\sigma$, then consider separately two cases: $c<50C\ell/\ln n$ and $c\geq 50C\ell/\ln n$. In the first case, we have that $\sigma\geq 2C\ell/\ln n$ due to (48). Therefore, for large enough $C$, due to (47),
$\Pi^{H}_{\ell}\left((1+o(1))\frac{N}{m^{\prime}}\right)^{\ell}$ ($\leq\frac{{\ell_{\chi}\choose c}2^{12\ell}n^{c/6}(e\ell_{\chi}/\sigma)^{\sigma}}{n^{\ell/2+c/2+\sigma/2}}\left((1+o(1))\frac{N}{m^{\prime}}\right)^{\ell}$
$\leq\left(\frac{3Cn^{1/6}}{c}\right)^{c}\left(\frac{3C}{\sigma}\right)^{\sigma}\left(\frac{2^{12}}{\varepsilon}+o(1)\right)^{\ell}<\left(\frac{n^{0.2}}{c}\right)^{c}e^{-C\ell/2}.$
In the latter case, since $c=\Omega(\ell/\ln n)$ is large, we get the bound
($\frac{{\ell_{\chi}\choose c}2^{12\ell}n^{c/6}(e\ell_{\chi}/\sigma)^{\sigma}}{n^{\ell/2+c/2+\sigma/2}}\left((1+o(1))\frac{N}{m^{\prime}}\right)^{\ell}$ $\leq\left(\frac{3Cn^{1/6}}{c}\right)^{c}\left(\frac{3C}{\sigma}\right)^{\sigma}\left(\frac{2^{12}}{\varepsilon}+o(1)\right)^{\ell}$
$<n^{-0.1c}\left(\frac{2^{12}}{\varepsilon}+o(1)\right)^{\ell}<e^{-C\ell}.$
Finally, let $n^{c/6}{\ell_{\chi}\choose\min\{\lfloor\ell_{\chi}/2\rfloor,\sigma\}}<{\ell_{\chi}\choose\min\{\lfloor\ell_{\chi}/2\rfloor,c+\sigma\}}$. If $\lfloor\ell_{\chi}/2\rfloor\leq c+\sigma$, we get exactly the same bound as in (50). We then assume that $\lfloor\ell_{\chi}/2\rfloor>c+\sigma$. With this assumption, we get that (${\ell_{\chi}\choose c}\leq{\ell_{\chi}\choose c+\sigma}$. Therefore,
$\Pi^{H}_{\ell}\left((1+o(1))\frac{N}{m^{\prime}}\right)^{\ell}$ ($\leq\frac{{\ell_{\chi}\choose c}2^{12\ell}{\ell_{\chi}\choose c+\sigma}}{n^{\ell/2+c/2+\sigma/2}}\left((1+o(1))\frac{N}{m^{\prime}}\right)^{\ell}\leq\frac{\left(\frac{e\ell_{\chi}}{c+\sigma}\right)^{2(c+\sigma)}(2^{12}/\varepsilon+o(1))^{\ell}}{n^{c/2+\sigma/2}}$
(48)${{\leq}}\left(\frac{9C^{2}n^{1/2}}{(c+\sigma)^{2}}\right)^{c+\sigma}\left(\frac{2^{12}}{\varepsilon}+o(1)\right)^{\frac{\sigma\ln n}{4C}+\frac{c\ln n}{100C}}<\left(\frac{n^{0.55}}{(c+\sigma)^{2}}\right)^{c+\sigma}.$